OpenStudy (anonymous):

if p varies inversely as q and p=4 when q=6, then which if the following represents another possible solution for p and q?

OpenStudy (anonymous):

p = k/q so 4 = k/6 giving k = 24 what follows?

OpenStudy (anonymous):

oops sorry... a)p=8 and q= 12 b.p=8 and q=10 c.p=10 and q=12 d.p=12 and q=1 e.p=12 and q=2

OpenStudy (anonymous):

p = 24/q so if p = 12 q = 2 (12 = 24/2) answer is e

OpenStudy (anonymous):

thank u soo much...:)